I need help with this problem. Thank you.
Determine the linear function whose graph is a line that parallel to the line g(x)=-3x-4 and contains the point (2,7).

Answer by scott8148(6628)

parallel means it has the same slope ... using point-slope gives (y-7)=-3(x-2)

y-7=-3x+6 ... y=-3x+13

f(x)=-3x+13